Circular No. 7556

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN PUPPIS I. Platais, V. Kozhurina-Platais, and M. I. Zacharias, Universities

Space Research Association; and N. Zacharias, U.S. Naval Observatory (USNO), report the following position for this variable star (cf. IAUC 7554) from preliminary reduction of two CCD frames (scale 0′′.905/pixel) obtained on Jan. 4.227 UT with the USNO 0.20-m f/10 astrograph at Cerro Tololo Interamerican Observatory: α = 7^{h}37^{m}56^{s}.882, δ = −25^{o}56′58′′.88 (equinox 2000.0; 100 Tycho-2 reference stars; estimated standard error 0′′.05 in each coordinate). The star’s red magnitude was measured to be 8.7. There is a star (red mag 13.6; epoch 1999 Mar. 3) in the first USNO CCD Astrographic Catalog (UCAC1) at position end figures 56^{s}.883, 58′′.88; the same star is also visible (visual mag 13.1) on a pair of the Yale/San Juan Southern Proper Motion program plates taken on 1967 Jan. 18. Preliminary reductions indicate that its absolute proper motion is small: μ_{α} cos δ = −0′′.0047/yr and μ_{δ} = +0′′.0064/yr (standard error 0′′.004/yr).

R. M. Wagner, LBT Observatory; C. B. Foltz, MMT Observatory; and S. G. Starrfield, Arizona State University, write: “A spectrum (range 420–750 nm, resolution 0.35 nm) of the possible nova, obtained on Jan. 4.23 with the 6.5-m MMT (+ blue-channel spectrograph), is dominated by emission lines arising from Fe II (multiplets 27, 37, 38, 42, 46, 48, 49, 55, 73, and 74); Na I 589.4 (1) and 616 (5) nm; and Ca I 658 (1) nm. The emission lines exhibit P-Cyg profiles, and overlapping emission and absorption makes positive spectral identifications difficult. Emission lines arising from H I and He I do not appear to be present. Measurement of several absorption troughs gives a maximum expansion velocity of ∼ 900 km/s. The spectrum is not typical of ‘Fe II’ classical novae, recurrent novae, or symbiotic novae such as PU Vul or RR Tel.”

Spectroscopic reports have also been received from several observers using smaller instruments: M. Fujii, Okayama, Japan (0.28-m reflector, 2000 Dec. 31.7 UT); W. Liller, Viña del Mar, Chile (0.20-m Schmidt telescope + objective prism, 2001 Jan. 3.235); and O. Shemmer, E. Leibowitz and J. Dan, Wise Observatory (Jan. 4.00). They all report the prominent emission lines of Fe II; though they also report He I at 587.6 nm, Wagner notes this could also be due to Na I 589.4-nm. Fujii’s spectrum (resolution ∼ 1 nm) suggests that Hβ is weak or absent; emission in the other spectra close to Hα is likely due to Ca I 658.0-nm, according to Wagner. Shemmer adds that the continuum rises by a factor of 1.5 from 750 to 400 nm, and the spectrum shows emission line widths of Fe II, O I, Na I, and Ca II that correspond to an expansion velocity of ∼ 1000 km/s.

2001 January 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green